OpenStudy (ashleyisakitty):

Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 12 feet by 12 feet by 9 inches for a patio if the concrete costs $54 per cubic yard?

OpenStudy (anonymous):

L*W*H 54 / volume of slab

OpenStudy (whpalmer4):

@DSS that won't get you the right answer

OpenStudy (ashleyisakitty):

I didnt think so. The answers I was getting with that were not logical.

OpenStudy (pratyush5):

remember that the units are different. You gotta interchange into either feet or inches

OpenStudy (whpalmer4):

You need to find the volume of the slab. Notice that you have some unit conversions to do. Then multiply the volume of the slab by the unit price of the concrete (again, using compatible units)

OpenStudy (pratyush5):

@whpalmer4 is right. Just dont mess up with the units.

OpenStudy (ashleyisakitty):

I dont think I did it right. (12/4)*(12/4)*(1/4)=9/4 (9/4)*54=121.5 121.5 is none of my answer choices.

OpenStudy (pratyush5):

wait let me do it .

OpenStudy (pratyush5):

is 1944 any of your options ?

OpenStudy (ashleyisakitty):

Yes that is an option!

OpenStudy (pratyush5):

BAZINGA :P

OpenStudy (ashleyisakitty):

Thanks :DD What did you differently than me? :x

OpenStudy (pratyush5):

First i calculated the volume in feet. Can you do it ?

OpenStudy (pratyush5):

calculate and tell me ?

OpenStudy (ashleyisakitty):

No, i dont think i understood how to do that. I just calculated how many times 12 would go into a yard.

OpenStudy (pratyush5):